If the sum of a number and 1 is tripled , the result is three less than twice the number . Find the number
Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(69443)

If the sum of a number and 1 is tripled , the result is three less than twice the number . Find the number
---------
n = the number
n+1 = the number and 1
------
3(n+1) = ... tripled
Can you do the rest?

Answer by lynnlo(4176)

3(x+1)=2x-3
3x+3=2x-3
3x-2x+3=2x-2x-3
x+3=-3
x+3-3=-3-3
x=-6